# CAX: Cellular Automata Accelerated in JAX

### v1

@@ Abstract
Cellular automata have become a cornerstone for investigating emergence and self-organization across diverse scientific disciplines, spanning neuroscience, artificial life, and theoretical physics. However, the absence of a hardware-accelerated cellular automata library limits the exploration of new research directions, hinders collaboration, and impedes reproducibility. In this work, we introduce CAX (Cellular Automata Accelerated in JAX), a high-performance and flexible open-source library designed to accelerate cellular automata research. CAX offers cutting-edge performance and a modular design through a user-friendly interface, and can support both discrete and continuous cellular automata with any number of dimensions. We demonstrate CAX’s performance and flexibility through a wide range of benchmarks and applications. From classic models like elementary cellular automata and Conway’s Game of Life to advanced applications such as growing neural cellular automata and self-classifying MNIST digits, CAX speeds up simulations up to 2,000 times faster. Furthermore, we demonstrate CAX’s potential to accelerate research by presenting a collection of three novel cellular automata experiments, each implemented in just a few lines of code thanks to the library’s modular architecture. Notably, we show that a simple one-dimensional cellular automaton can outperform GPT-4 on the 1D-ARC challenge.
@@ 1 Introduction
In recent years, the integration of machine learning techniques with cellular automata has opened new avenues for research in morphogenesis (Mordvintsev et al., 2020), self-organization (Randazzo et al., 2020; 2021), and developmental processes (Najarro et al., 2022). The advent of Neural Cellular Automata (NCA) has significantly broadened the scope of CA research, yielding profound biological insights and showcasing the power of gradient-based optimization in studying emergence and self-organization. NCAs extend traditional CAs by incorporating neural networks to learn update rules, allowing for more complex and adaptive behaviors. This approach enables the modeling of sophisticated phenomena such as pattern formation in biological systems and the evolution of artificial life forms. This progress has not only deepened our understanding of complex systems but also underscored the growing computational demands of CA experiments, pointing to the potential for scaling through hardware-accelerated libraries inspired by advances in deep learning.
@@
In response to these challenges and opportunities, we present CAX: Cellular Automata Accelerated in JAX, an open-source library with cutting-edge performance, designed to provide a flexible and efficient framework for cellular automata research. CAX is built on JAX (Bradbury et al., 2018), a high-performance numerical computing library, enabling to speed up cellular automata simulations through massive parallelization across various hardware accelerators such as CPUs, GPUs, and TPUs. CAX is flexible and supports both discrete and continuous cellular automata with any number of dimensions, accommodating classic models like elementary cellular automata and Conway’s Game of Life, as well as modern variants such as Lenia and Neural Cellular Automata.
@@
CAX’s flexibility and potential to accelerate research is showcased through three novel cellular automata experiments. Thanks to CAX’s modular architecture, each of these experiments is implemented in just a few lines of code (Appendix B), significantly reducing the barrier to entry for cellular automata research. Notably, we show that a simple one-dimensional cellular automaton implemented with CAX outperforms GPT-4 on the 1D-ARC challenge (Xu et al., 2024), see Section 5.3. Finally, to support users and facilitate adoption, CAX comes with high-quality, diverse examples and comprehensive documentation. The list of implemented CAs is detailed in Table 1.
@@ 2.1 Cellular Automata
$\Phi(\mathbf{S}_{0})=\mathbf{S}_{1},\Phi(\mathbf{S}_{1})=\mathbf{S}_{2},\dots$
@@ 2.2 Controllable Cellular Automata
$\Phi(\mathbf{S}_{0},\mathbf{I}_{0})=\mathbf{S}_{1},\Phi(\mathbf{S}_{1},\mathbf{I}_{1})=\mathbf{S}_{2},\dots$
@@ 3 CAX: Cellular Automata Accelerated in JAX
Figure 2: High-level architecture of CAX, illustrating the modular design with perceive and update components. This flexible structure supports various CA types across multiple dimensions. (Adapted from Mordvintsev et al. (2020) under CC-BY 4.0 license.)
CAX is a high-performance and flexible open-source library designed to accelerate cellular automata research. In this section, we detail CAX’s architecture, design and key features. At its core, CAX leverages JAX and Flax (Heek et al., 2024), capitalizing on the well-established connection between CA and recurrent convolutional neural networks. This synergy, discussed in Section 2), allows CAX to harness advancements in machine learning to accelerate CA research. CAX offers a modular and intuitive design through a user-friendly interface, supporting both discrete and continuous cellular automata across any number of dimensions. This flexibility enables researchers to seamlessly transition between different CA types and complexities within a single, unified framework (Table 1). We have made our anonymized repository available at github.com/b769eb6f/cax. We invite readers to experience CAX’s capabilities firsthand by accessing our curated examples as interactive notebooks in Google Colab, conveniently linked in the repository’s README.
@@ 3.1 Architecture and Design
CAX introduces a unifying framework for all cellular automata types, encompassing discrete, continuous, and neural models across any number of dimensions (Table 1). This flexible architecture is built upon two key components: the perceive module and the update module. Together, these modules define the local rule of the CA. At each time step, this local rule is applied uniformly to all cells in the grid, generating the next global state of the system, as explained in Section 2.1. This modular approach not only provides a clear separation of concerns but also facilitates easy experimentation and extension of existing CA models.
@nnx.jit
def step(self, state: State, input: Input | None = None) -> State:
"""Perform a single step of the CA.
Args:
state: Current state.
input: Optional input.
Returns:
Updated state.
perception = self.perceive(state)
state = self.update(state, perception, input)
return state
@@ 3.2.1 Performance
def step(carry: tuple[CA, State], input: Input | None) -> tuple[tuple[CA, State], State]:
ca, state = carry
state = ca.step(state, input)
return (ca, state), state if all_steps else None
(_, state), states = nnx.scan(
step,
in_axes=(nnx.Carry, input_in_axis),
length=num_steps,
)((self, state), input)
@@ 3.2.2 Utilities
CAX offers a rich set of utility functions to support various aspects of cellular automata research. A high-quality implementation of the sampling pool technique is provided, which is crucial for training stable growing neural cellular automata Mordvintsev et al. (2020). To facilitate the training of unsupervised neural cellular automata and enable generative modeling within the CA framework, CAX incorporates a variational autoencoder implementation. Additionally, the library provides utilities for handling image and emoji inputs, allowing for diverse and visually engaging CA experiments. These utilities are designed to streamline common tasks in CA research, allowing researchers to focus on their specific experiments rather than reimplementing standard components.
@@ 5.1 Diffusing Neural Cellular Automata
The conventional NCA training method typically employs a ”sample pool” strategy to address stability issues and encourage the formation of attractors. This approach involves maintaining a diverse pool of intermediate states, sampling from this pool for training, and periodically updating it with newly generated states. By exposing the NCA to various intermediate configurations and consistently guiding them towards the target pattern, the sample pool method helps shape the system’s dynamics, making the desired pattern a more robust attractor in the state space.
Our proposed diffusion-inspired approach offers several advantages over the traditional growing mechanism. First, unlike the growing mechanism, our diffusion-based approach doesn’t require a sample pool, which simplifies the training process and reduces memory requirements, making it more efficient and scalable. Second, our diffusion-inspired approach naturally guides the NCA towards more stable dynamics, effectively creating a stronger attractor basin around the target pattern. In Figure 5, we compare the regeneration capabilities of growing NCAs with diffusing NCAs. We create an artificial damage by cutting the tail of the gecko and observe that diffusing NCA demonstrate emergent regenerating capabilities.
@@ 5.2 Self-autoencoding MNIST Digits
In this experiment, we draw inspiration from Randazzo et al. (2020) where a NCA is trained to classify MNIST digits through local interactions. In their work, each cell (pixel) of an MNIST digit learns to output the correct digit label through local communication with neighboring cells. The NCA demonstrates the ability to reach global consensus on digit classification, maintain this classification over time, and adapt to perturbations or mutations of the digit shape. Their model showcases emergent behavior, where simple local rules lead to complex global patterns, analogous to biological systems achieving anatomical homeostasis.
@@
In this experiment, we train a one-dimensional NCA on the 1D-ARC dataset (Xu et al., 2024). The 1D-ARC dataset is a novel adaptation of the original Abstraction and Reasoning Corpus (Chollet, 2019) (ARC), designed to simplify and streamline research in artificial intelligence and language models. By reducing the dimensionality of input and output images to a single row of pixels, 1D-ARC maintains the core knowledge priors of ARC while significantly reducing task complexity. For example, the tasks in 1D-ARC include ”Static movement by 3 pixels”, ”Fill”, and ”Recolor by Size Comparison”. For a full description of the dataset, see the project page.
@@
The primary goal of this experiment is for the NCA to learn from the training set, a generalizable rule for each task that can solve unseen examples from the test sets. This challenge tests the NCA’s ability to infer abstract patterns and apply them to new situations, a key aspect of human-like reasoning. Figure 8 illustrates the NCA’s “reasoning” on all 1D-ARC tasks. The visualization shows the input at the top, intermediate steps, and final output of the NCA at the bottom of each image, and is called a space-time diagram.
To evaluate the NCA’s performance, we compare it to GPT-4, a state-of-the-art language model, on the 1D-ARC test set. Table 2 presents the accuracy of the NCA and GPT-4 across 18 different task types. The GPT-4 values are direct-grid results, directly taken from Xu et al. (2024) Appendix A. Notably, the NCA outperforms GPT-4 on several tasks, particularly those involving movement, pattern copying, and denoising. Overall, the NCA achieves a total accuracy of 60.12% compared to GPT-4’s 41.56%, as reported by Xu et al. (2024).
These results demonstrate the potential of NCAs in solving abstract reasoning tasks, even outperforming sophisticated language models in certain domains. The NCA’s success in tasks like ”Move 3” and ”Pattern Copy Multicolor” showcases its ability to learn complex spatial transformations and apply them consistently.
@@ Appendix A Hyperparameters
In this section, we detail the hyperparameters for the three novel neural cellular automata experiments presented in Section 5.
Table 3: Diffusing Neural Cellular Automata
@@
Table 4: Self-autoencoding MNIST Digits
@@
Table 5: 1D-ARC Neural Cellular Automata
@@
Appendix B Example Notebook
## Import
import jax
import jax.numpy as jnp
import mediapy
import optax
from cax.core.ca import CA
from cax.core.perceive.depthwise_conv_perceive import DepthwiseConvPerceive
from cax.core.perceive.kernels import grad_kernel, identity_kernel
from cax.core.state import state_from_rgba_to_rgb, state_to_rgba
from cax.core.update.nca_update import NCAUpdate
from cax.nn.pool import Pool
from cax.utils.image import get_emoji
from flax import nnx
from tqdm.auto import tqdm
## Configuration
seed = 0
channel_size = 16
num_kernels = 3
hidden_size = 128
cell_dropout_rate = 0.5
pool_size = 1_024
batch_size = 8
num_steps = 128
learning_rate = 2e-3
emoji = "gecko"
target_size = 40
target_padding = 16
key = jax.random.key(seed)
rngs = nnx.Rngs(seed)
## Dataset
target = get_emoji(emoji, size=target_size, padding=target_padding)
## Init state
def init_state():
state_shape = target.shape[:2] + (channel_size,)
state = jnp.zeros(state_shape)
mid = tuple(size // 2 for size in state_shape[:-1])
return state.at[mid[0], mid[1], -1].set(1.0)
## Model
perceive = DepthwiseConvPerceive(channel_size, rngs)
update = NCAUpdate(channel_size, num_kernels * channel_size, (hidden_size,), rngs, cell_dropout_rate=cell_dropout_rate)
kernel = jnp.concatenate([identity_kernel(ndim=2), grad_kernel(ndim=2)], axis=-1)
kernel = jnp.expand_dims(jnp.concatenate([kernel] * channel_size, axis=-1), axis=-2)
perceive.depthwise_conv.kernel = nnx.Param(kernel)
ca = CA(perceive, update)
## Train
state = jax.vmap(lambda _: init_state())(jnp.zeros(pool_size))
pool = Pool.create({"state": state})
lr_sched = optax.linear_schedule(init_value=learning_rate, end_value=0.1 * learning_rate, transition_steps=2_000)
optimizer = optax.chain(
optax.clip_by_global_norm(1.0),
optax.adam(learning_rate=lr_sched),
update_params = nnx.All(nnx.Param, nnx.PathContains("update"))
optimizer = nnx.Optimizer(ca, optimizer, wrt=update_params)
def mse(state):
return jnp.mean(jnp.square(state_to_rgba(state) - target))
@nnx.jit
def loss_fn(ca, state, key):
state_axes = nnx.StateAxes({nnx.RngState: 0, ...: None})
state = nnx.split_rngs(splits=batch_size)(
nnx.vmap(
lambda ca, state: ca(state, num_steps=num_steps, all_steps=True),
in_axes=(state_axes, 0),
)(ca, state)
# Sample a random step
index = jax.random.randint(key, (state.shape[0],), num_steps // 2, num_steps)
state = state[jnp.arange(state.shape[0]), index]
loss = mse(state)
return loss, state
@nnx.jit
def train_step(ca, optimizer, pool, key):
sample_key, loss_key = jax.random.split(key)
# Sample from pool
pool_index, batch = pool.sample(sample_key, batch_size=batch_size)
current_state = batch["state"]
# Sort by descending loss
sort_index = jnp.argsort(jax.vmap(mse)(current_state), descending=True)
pool_index = pool_index[sort_index]
current_state = current_state[sort_index]
# Sample a new target to replace the worst
new_state = init_state()
current_state = current_state.at[0].set(new_state)
(loss, current_state), grad = nnx.value_and_grad(loss_fn, has_aux=True, argnums=nnx.DiffState(0, update_params))(
ca, current_state, loss_key
optimizer.update(grad)
pool = pool.update(pool_index, {"state": current_state})
return loss, pool
num_train_steps = 8_192
print_interval = 128
pbar = tqdm(range(num_train_steps), desc="Training", unit="train_step")
losses = []
for i in pbar:
key, subkey = jax.random.split(key)
loss, pool = train_step(ca, optimizer, pool, subkey)
losses.append(loss)
if i % print_interval == 0 or i == num_train_steps - 1:
avg_loss = sum(losses[-print_interval:]) / len(losses[-print_interval:])
pbar.set_postfix({"Average Loss": f"{avg_loss:.6f}"})

### v2

@@ Abstract
Cellular automata have become a cornerstone for investigating emergence and self-organization across diverse scientific disciplines. However, the absence of a hardware-accelerated cellular automata library limits the exploration of new research directions, hinders collaboration, and impedes reproducibility. In this work, we introduce CAX (Cellular Automata Accelerated in JAX), a high-performance and flexible open-source library designed to accelerate cellular automata research. CAX delivers cutting-edge performance through hardware acceleration while maintaining flexibility through its modular architecture, intuitive API, and support for both discrete and continuous cellular automata in arbitrary dimensions. We demonstrate CAX’s performance and flexibility through a wide range of benchmarks and applications. From classic models like elementary cellular automata and Conway’s Game of Life to advanced applications such as growing neural cellular automata and self-classifying MNIST digits, CAX speeds up simulations up to 2,000 times faster. Furthermore, we demonstrate CAX’s potential to accelerate research by presenting a collection of three novel cellular automata experiments, each implemented in just a few lines of code thanks to the library’s modular architecture. Notably, we show that a simple one-dimensional cellular automaton can outperform GPT-4 on the 1D-ARC challenge.
@@ 1 Introduction
In recent years, the integration of machine learning techniques with cellular automata has opened new avenues for research in morphogenesis (Mordvintsev et al., 2020) and self-organization (Randazzo et al., 2020; 2021; Najarro et al., 2022). Neural Cellular Automata (NCA) represent a key advancement in this field, using differentiable neural networks to learn update rules through gradient-based optimization rather than relying on manual design. Recent works have combined NCA with advanced architectures like Convolutional Neural Networks (Pajouheshgar et al., 2023), Graph Neural Networks (Grattarola et al., 2021) and Vision Transformers (Tesfaldet et al., 2022). The convergence of machine learning and cellular automata research not only deepened our understanding of complex systems but also underscored the growing computational demands of CA experiments.
@@
In response to these challenges and opportunities, we present CAX: Cellular Automata Accelerated in JAX, an open-source library with cutting-edge performance, designed to provide a flexible and efficient framework for cellular automata research. CAX is built on JAX (Bradbury et al., 2018), a high-performance numerical computing library, enabling to speed up cellular automata simulations through massive parallelization across various hardware accelerators such as CPUs, GPUs, and TPUs. CAX is flexible and supports both discrete and continuous cellular automata with any number of dimensions, accommodating classic models like elementary cellular automata and Conway’s Game of Life, as well as modern variants such as Lenia and Neural Cellular Automata (Figure 1).
@@
CAX’s flexibility and potential to accelerate research is showcased through three novel cellular automata experiments. Thanks to CAX’s modular architecture, each of these experiments is implemented in just a few lines of code, significantly reducing the barrier to entry for cellular automata research. Notably, we show that a simple one-dimensional cellular automaton implemented with CAX outperforms GPT-4 on the 1D-ARC challenge (Xu et al., 2024), see Section 5.3. Finally, to support users and facilitate adoption, CAX comes with high-quality, diverse examples and comprehensive documentation. The list of implemented CAs is detailed in Table 1.
@@ 2.1 Cellular Automata
$\mathbf{S}_{1}=\Phi(\mathbf{S}_{0}),\mathbf{S}_{2}=\Phi(\mathbf{S}_{1}),\dots$
@@ 2.2 Controllable Cellular Automata
$\mathbf{S}_{1}=\Phi(\mathbf{S}_{0},\mathbf{I}_{0}),\mathbf{S}_{2}=\Phi(\mathbf{S}_{1},\mathbf{I}_{1}),\dots$
@@ 3 CAX: Cellular Automata Accelerated in JAX
Figure 2: High-level architecture of CAX, illustrating the modular design with perceive and update modules. This flexible structure supports various CA types across multiple dimensions. (Adapted from Mordvintsev et al. (2020) under CC-BY 4.0 license.)
CAX is a high-performance and flexible open-source library designed to accelerate cellular automata research. In this section, we detail CAX’s architecture, design and key features. At its core, CAX leverages JAX and Flax (Heek et al., 2024), capitalizing on the well-established connection between CA and recurrent convolutional neural networks. This synergy, discussed in Section 2, allows CAX to harness advancements in machine learning to accelerate CA research. CAX offers a modular and intuitive design through a user-friendly interface, supporting both discrete and continuous cellular automata across any number of dimensions. This flexibility enables researchers to seamlessly transition between different CA within a single, unified framework (Table 1). We invite readers to experience CAX’s capabilities firsthand by accessing our examples as interactive notebooks in Google Colab, conveniently linked in the README of the repository https://github.com/maxencefaldor/cax.
@@ 3.1 Architecture and Design
CAX introduces a unifying framework for all cellular automata types. This flexible architecture is built upon two key components: the perceive module and the update module, see Figure 2. Together, these modules define the local rule of the CA. At each time step, this local rule is applied uniformly to all cells in the grid, generating the next global state of the system, as explained in Section 2.1. This modular approach not only provides a clear separation of concerns but also facilitates easy experimentation and extension of existing CA models.
Building on the controllable CA framework introduced in Section 2.2, our architecture generalizes the neural cellular automata approach to recurrent convolutional neural architectures, enabling seamless integration of external inputs and control signals. By implementing both modules using standard machine learning components from Flax, CAX makes it straightforward to experiment with various neural architectures while maintaining the cellular automata paradigm - from simple convolutional layers to sophisticated attention mechanisms.
@@ 3.2.2 Utilities
CAX offers a rich set of utility functions to support various aspects of cellular automata research. A high-quality implementation of the sampling pool technique is provided, which is crucial for training stable growing neural cellular automata (Mordvintsev et al., 2020). To facilitate the training of unsupervised neural cellular automata and enable generative modeling within the CA framework, CAX incorporates a variational autoencoder implementation. Additionally, the library provides utilities for handling image and emoji inputs, allowing for diverse and visually engaging CA experiments. These utilities are designed to streamline common tasks in CA research, allowing researchers to focus on their specific experiments rather than reimplementing standard components.
@@ 5.1 Diffusing Neural Cellular Automata
The conventional NCA training method typically employs a “sample pool” strategy to address stability issues and encourage the formation of attractors. This approach involves maintaining a diverse pool of intermediate states, sampling from this pool for training, and periodically updating it with newly generated states. By exposing the NCA to various intermediate configurations and consistently guiding them towards the target pattern, the sample pool method helps shape the system’s dynamics, making the desired pattern a more robust attractor in the state space.
Our proposed diffusion-inspired approach offers several advantages over the traditional growing mechanism. First, unlike the growing mechanism, our diffusion-based approach does not require a sample pool, which simplifies the training process and reduces memory requirements, making it more efficient and scalable. Second, our diffusion-inspired approach naturally guides the NCA towards more stable dynamics, effectively creating a stronger attractor basin around the target pattern. In Figure 5, we compare the regeneration capabilities of growing NCAs with diffusing NCAs. We create an artificial damage by cutting the tail of the gecko and observe that diffusing NCA demonstrate emergent regenerating capabilities.
@@ 5.2 Self-autoencoding MNIST Digits
In this experiment, we draw inspiration from Randazzo et al. (2020) where an NCA is trained to classify MNIST digits through local interactions. In their work, each cell (pixel) of an MNIST digit learns to output the correct digit label through local communication with neighboring cells. The NCA demonstrates the ability to reach global consensus on digit classification, maintain this classification over time, and adapt to perturbations or mutations of the digit shape. Their model showcases emergent behavior, where simple local rules lead to complex global patterns, analogous to biological systems achieving anatomical homeostasis.
@@
In this experiment, we train a one-dimensional NCA on the 1D-ARC dataset (Xu et al., 2024). The 1D-ARC dataset is a novel adaptation of the original Abstraction and Reasoning Corpus (Chollet, 2019) (ARC), designed to simplify and streamline research in artificial intelligence and language models. By reducing the dimensionality of input and output images to a single row of pixels, 1D-ARC maintains the core knowledge priors of ARC while significantly reducing task complexity. For example, the tasks in 1D-ARC include “Static movement by 3 pixels”, “Fill”, and “Recolor by Size Comparison”. For a full description of the dataset, see the project page.
@@
The primary goal of this experiment is for the NCA to learn a generalizable rule from the training set, enabling it to solve unseen examples in the test sets. This challenge tests the NCA’s ability to infer abstract patterns and apply them to new situations, a key aspect of human-like reasoning. Figure 8 illustrates the NCA’s “reasoning” on all 1D-ARC tasks. The visualization shows the input at the top, intermediate steps, and final output of the NCA at the bottom of each image, and is called a space-time diagram.
To evaluate the NCA’s performance, we compare it to GPT-4, a state-of-the-art language model, on the 1D-ARC test set. Table 2 presents the accuracy of the NCA and GPT-4 across 18 different task types. The GPT-4 values are direct-grid results, directly taken from Xu et al. (2024). Notably, the NCA outperforms GPT-4 on several tasks, particularly those involving movement, pattern copying, and denoising. Overall, the NCA achieves a total accuracy of 60.12% compared to GPT-4’s 41.56%, as reported by Xu et al. (2024).
These results demonstrate the potential of NCAs in solving abstract reasoning tasks, even outperforming sophisticated language models in certain domains. The NCA’s success in tasks like “Move 3” and “Pattern Copy Multicolor” showcases its ability to learn complex spatial transformations and apply them consistently.
@@ 6 Conclusion
Acknowledgments
We thank Gabriel Béna and Arthur Braida for insightful discussions and feedback. We also acknowledge the financial support provided by a grant from G-Research.
@@ Appendix A Hyperparameters
This appendix provides detailed hyperparameters for the three novel neural cellular automata (NCA) experiments introduced in Section 5. These hyperparameters govern the architecture, training process, and simulation dynamics of each experiment. For further details on the experimental setup, refer to the respective subsections in Section 5 or to the notebooks on https://github.com/maxencefaldor/cax.
Table 3: Hyperparameters for Diffusing Neural Cellular Automata (see Section 5.1)
@@
Table 4: Hyperparameters for Self-autoencoding MNIST Digits (see Section 5.2)
@@
Table 5: Hyperparameters for 1D-ARC Neural Cellular Automata (see Section 5.3)